# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ B.2 Pre-processing
Unable to obtain SELFIES. We retrieve the SMILES representation of a molecule with its CID using pubchempy (Swain, 2024) API, for CIDs that cannot be found with pubchempy.Compound.from_cid(), we discard them. For molecules that cannot be converted to SELFIES, we discard them. This could happen when processing datasets from Li et al. (2024b).
Overlapped samples. Datasets from different sources often contain overlapping samples, leading to potential data leakage. For example, solvent and catalyst prediction are subsets of reagent prediction, and molecule description data from 3D-MOIT (Li et al., 2024b) may include samples that overlap with those in ChEBI-20 (Edwards et al., 2021). Such overlaps create scenarios where a sample from one dataset’s training set appears in the test set of another, compromising the reliability of model evaluation. To address this issue, we conduct a thorough dataset comparison to identify potential overlaps and systematically remove any samples from the training sets that also appear in the test sets of other datasets.
B.3 Evaluation Metrics
@@
$\mathbf{R}^{2}$. The $R^{2}$ metric scores the proportion of variability in the target variable that can be explained by the model’s predictors. It can be served as an indicator of the model’s explanatory strength, reflecting how well the observed data points are captured by the regression model.
@@ C.1 Graph Tokenizer
Molecule to graph conversion. Graph neural network is widely used in many scenarios like traffic (Zhang et al., 2025), society relationships (Zhang et al., 2024a), and also molecules (Sun et al., 2022, 2024). Following vanilla setting, we utilize RDKit (Landrum et al., 2013) to transform SELFIES into graph structure in our experiments. For tasks involving a single molecule as input, the molecule is converted directly. For tasks requiring multiple molecules as input, only the first molecule in the input sequence is converted into a graph. Our model does not incorporate multi-graph understanding; instead, it processes both the graph and SELFIES representation of the first molecule, while only the SELFIES representations are provided for the remaining molecules. Meanwhile, since MoleculeSTM (Liu et al., 2023) incorporates additional molecular graph-text contrastive training compared to GraphMVP (Liu et al., 2022), leading to improved multimodal model training efficiency, we adopt MoleculeSTM as the graph encoder.
@@ C.2 Multimodal Alignment
To balance the molecular graph and text modalities while ensuring training efficiency, we employ a single-layer linear projector in Stage 1. Following Liu et al. (2023), we carefully filter PubChem to obtain 310K+ graph-text pairs and convert them into instruction-following data for pretraining. The alignment between the molecular graph and text modalities is enhanced solely by adjusting the parameters of the single-layer linear projector.
C.3 Parameter-efficient Fine-tuning (PEFT)
As the size of recent models increases rapidly, updating the models in parameter-efficient ways becomes crucial. PEFT (Ding et al., 2023; Wang et al., 2023; Zhai et al., 2023; Yu et al., 2024a; Wang et al., 2024) methods diverge from the conventional approach of fine-tuning the entire pre-trained model, instead only learning a few additional parameters for knowledge transfer. Due to the redundancies of attention matrix in LLMs’ pre-trained parameter $\mathcal{W}_{0}$, we hope to implement the low-rank approximation (LoRA) of the tensor $\mathcal{W}_{0}$ to get the new learnable weight tensor $\Delta\mathcal{W}\in\mathbb{R}^{m\times n}$ for downstream knowledge transfer. $\Delta\mathcal{W}$ can be constructed with the product of two lower-dimensional matrices $\mathcal{A}\in\mathbb{R}^{m\times r}$ and $\mathcal{B}\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$. The goal is to find $\mathcal{A}$ and $\mathcal{B}$ that minimize the approximation error between $W$ and $\mathcal{A}\mathcal{B}$. A common objective is the Frobenius norm, leading to the minimization problem $\min_{\mathcal{A},\mathcal{B}}\,\|\Delta\mathcal{W}-\mathcal{A}\mathcal{B}\|_{F}^{2}$. Here, $\mathcal{A}$ and $\mathcal{B}$ together have significantly fewer parameters than $\Delta\mathcal{W}$ itself, making $\mathcal{A}\mathcal{B}$ an effective rank-$r$ approximation. Therefore, the backward propagation on the downstream fine-tuning data $\mathcal{D}$ can be expressed as:
$\nabla_{\boldsymbol{\Delta{\mathcal{W}}}}=\frac{\partial\mathcal{L}\left(\mathcal{D};\boldsymbol{\mathcal{W}}_{0}+\boldsymbol{\Delta{\mathcal{W}}}\right)}{\partial\boldsymbol{\Delta{\mathcal{W}}}}$ (11)
However, in Figure 2, we observe that the conventional LoRA approach can encounter intricate divergences in the middle to later phases of training, primarily due to a gradual increase in gradient norm values. Motivated by work (Kalajdzievski, 2023; Team, 2024), we attribute the main cause of these divergences to issues arising from the softmax operation when handling tasks drawn from different domains with substantially varied entropy, stemming from softmax’s translation invariance (i.e., $softmax\left(z\right)=softmax\left(z+c\right)$). Since all model parameters are shared among multiple tasks, each task competes by incrementally growing its norms. while this is not immediately detrimental, it leads to divergence once norms extend beyond the effective range of bf16. To mitigate this, we introduce the Adaptive Gradient Stabilization module in Section 3.4, employing an adaptive coefficient $\gamma_{\theta}=\frac{\alpha_{\theta}}{||r||_{p}}+\beta_{\theta}$, where $\alpha_{\theta}$ and $\beta_{\theta}$ are learnable variables and $r$ represents the rank. Through this mechanism, gradients can be adaptively stabilized in the form of:
$\nabla_{\boldsymbol{\Delta{\mathcal{W}}}}=\frac{\partial\mathcal{L}\left(\mathcal{D};\boldsymbol{\mathcal{W}}_{0}+\gamma_{\theta}\cdot\boldsymbol{\Delta{\mathcal{W}}}\right)}{\partial\boldsymbol{\Delta{\mathcal{W}}}}$ (12)
C.4 Frozen Pre-trained Backbone
We adopt LLaMA 3 (Dubey et al., 2024) as the backbone, a standard dense Transformer (Vaswani, 2017) architecture. It employs grouped query attention (Ainslie et al., 2023), which generalizes multi-query attention by introducing an intermediate set of key-value heads. Furthermore, LLaMA 3 applies an attention mask that blocks cross-document self-attention within a single sequence; while this feature shows minimal influence during standard pre-training, it becomes crucial for continued pre-training on long sequences. Lastly, LLaMA 3 supports an expanded vocabulary of 128K tokens.
C.5 Mutual Representation Similarity
Task scaling setup. We build a sequence of multi-task dataset with detailed composition as follows:
@@ C.5 Mutual Representation Similarity
  • 2 tasks: Reagent Prediction $+$ Molecular Captioning.
  • 4 tasks: Reagent Prediction $+$ Molecular Captioning $+$ Solvent Prediction $+$ Catalyst Prediction.
  • 8 tasks: Reagent Prediction $+$ Molecular Captioning $+$ Solvent Prediction $+$ Catalyst Prediction $+$ Forward Prediction $+$ Retrosynthesis $+$ Property Prediction $+$ Yield Regression.
@@
where $R^{\prime}\in\mathbb{R}^{B\times L\times d}$, and $m\in\mathbb{R}^{B\times T}$ is the mask indicating the padding tokens. We then flatten the first two dimensions and get $R^{\prime\prime}\in\mathbb{R}^{(B*L)\times d}$ and calculate the similarity with mutual_knn (Huh et al., 2024).
Let $N=B*L$, and we have two models $A$ and $B$ trained on different multi-task datasets, we first find their $k$ nearest neighbors $\text{knn}^{A}$ and $\text{knn}^{B}$.
@@ Appendix D Further Details on Training
Forward Reaction Prediction 8e-5 15 cosine 10 32 0.0075
@@
Omni-Molecular Tasks 15 15
@@
Specialist models are typically fine-tuned on a single task at a time, repeating the process separately for each task, a strategy known as separate tuning. In contrast, generalist models undergo simultaneous fine-tuning across multiple tasks, a process referred to as unified tuning. In this section, we present a detailed training framework for both of them on all experiments.
Separate instruction tuning. We follow the training recipe outlined in Cao et al. (2025). However, we observe significant overfitting when training the model on the molcap task for 20–50 epochs, as suggested in Cao et al. (2025). To address this issue, we manually allocate $10\%$ of the training set for validation and re-evaluated all tasks, we find that the recipes for forward prediction, reagent prediction, retrosynthesis, Quantum Mechanics Property Prediction from the original paper matches our results, however, we identify an updated training strategy tailored to the molcap task. The revised training recipe is summarized in Table 4.
@@
Environment Settings. We employ common techniques to boost performance and conserve memory, including FlashAttention 2 (Dao et al., 2022), activation checkpointing, and bf16 training. All experiments are conducted on 8×8$\times$8 ×NVIDIA A100 GPUs (80GB). For all experiments, the weight decay is set to 0. The term Stop Epoch in Table 4 shows the epoch that the experiment stops. This is because the early stop mechanism we used to prevent overfitting.
@@ Appendix E Further Details on Experimental Results
In this section, we provide a complete performance evaluation of Omni-Mol across the remaining 9 tasks.
Model Type #Param Weight(g/mol) LogP TPSA(Å2) Complexity SCF($10^{5}$eV)
Llama2 Specialist 6.7B 22.10 (96%) 1.45 (95%) 15.87 (92%) 69.74 (93%) 0.70 (99%)
3D-MoLM(S) Specialist 6.7B 14.79 (95%) 0.66 (97%) 9.71 (93%) 44.85 (94%) 0.35 (99%)
3D-MoLM(G) Generalist 6.7B 16.58 (92%) 0.78 (95%) 10.90 (90%) 45.49 (89%) 0.38 (98%)
Omni-Mol Generalist 1.7B 15.08 (100%) 0.59 (100%) 11.17 (100%) 49.38 (100%) 0.55 (99%)
Model Type #Param B-H S-M
Llama2 - 6.7B -0.476 0.121
Vicuna v1.5 - 6.7B -0.131 0.151
PRESTO Generalist 6.7B 0.944 0.652
Omni-Mol Generalist 1.7B 0.891 0.560
Table 5: More results for regression tasks. (Left) Results of property regression tasks, we report MAE for each task with the valid answer rate (%), since LMs sometimes fail to generate numerical responses corresponding to the given prompts. (Right) Results of Yield regression task, we report $R^{2}$ score.
@@
Model Type #Param Exact BLEU Levenshtein RDK MACCS Morgan Validity
Catalyst Prediction (PRESTO)
Llama2 - 6.7B 0.680 0.720 2.545 0.882 0.868 0.687 1.000
Vicuna v1.5 - 6.7B 0.685 0.703 2.451 0.883 0.869 0.692 1.000
nach0-base - - 0.000 0.072 36.442 0.129 0.055 0.009 0.849
Mol-Instruction Specialist 6.7B 0.000 0.110 28.424 0.031 0.045 0.015 0.999
T5Chem - - 0.022 0.346 13.408 0.146 0.268 0.200 0.996
PRESTO Generalist 6.7B 0.768 0.814 1.755 0.914 0.895 0.774 1.000
Omni-Mol Generalist 1.7B 0.752 0.860 1.544 0.919 0.903 0.759 1.000
Solvent Prediction (PRESTO)
Llama2 - 6.7B 0.311 0.462 3.819 0.452 0.480 0.417 1.000
Vicuna v1.5 - 6.7B 0.320 0.436 3.809 0.459 0.486 0.427 1.000
nach0-base - - 0.000 0.072 36.442 0.129 0.055 0.009 0.849
Mol-Instruction Specialist 6.7B 0.000 0.155 25.117 0.030 0.122 0.035 1.000
T5Chem - - 0.083 0.311 16.224 0.458 0.424 0.397 0.995
PRESTO Generalist 6.7B 0.419 0.695 2.758 0.529 0.547 0.506 0.912
Omni-Mol Generalist 1.7B 0.590 0.799 2.243 0.740 0.733 0.715 1.000
Table 6: More results for reaction tasks.
Model Type #Param BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
TextChemT5 Generalist 220M 0.541 0.406 0.615 0.403 0.564
MolT5-Large Specialist 780M 0.545 0.410 0.625 0.409 0.572
Galactica Specialist 1.3B 0.535 0.395 0.609 0.386 0.552
MolCA, Galac Specialist 1.3B 0.549 0.415 0.625 0.404 0.570
ReactXT, Galac Specialist 1.3B 0.574 0.440 0.644 0.427 0.589
Omni-Mol Generalist 1.7B 0.569 0.445 0.523 0.270 0.460
Table 7: Results of experimental procedure prediction.

### v3

@@ B.2 Preprocessing
Unable to obtain SELFIES. We retrieve the SMILES representation of a molecule with its CID using pubchempy [63] API, for CIDs that cannot be found with pubchempy.Compound.from_cid(), we discard them. For molecules that cannot be converted to SELFIES, we discard them.
Overlapped samples. Datasets from different sources often contain overlapping samples, leading to potential data leakage. For example, solvent and catalyst prediction are subsets of reagent prediction, and molecule description data from PubChem [33] may include samples that overlap with those in ChEBI-20 [20]. Such overlaps create scenarios where a sample from one dataset’s training set appears in the test set of another, compromising the reliability of model evaluation. To address this issue, we conduct a thorough dataset comparison to identify potential overlaps and systematically remove any samples from the training sets that also appear in the test sets of other datasets.
From SELFIES to 3D molecule. Compared to the 1D and 2D representations of molecules, the 3D spatial topology reveals richer molecular properties and research value, such as protein interactions and molecular dynamics. Following [41], we preprocess molecular SELFIES using RDKit. First, we convert SELFIES to a 2D representation via the selfies library and RDKit and add hydrogen atoms to facilitate subsequent force field optimization. Finally, after embedding each atom at a random initial coordinate, we optimize the resulting conformations using RDKit’s MMFF94 force field (e.g., by computing interatomic interaction potentials). Once we obtain the optimized atomic conformations, we execute 3D molecular tasks using the full 3D information.
B.3 Details on Evaluation Metrics
@@
$\mathbf{R}^{2}$. The $R^{2}$ metric scores the proportion of variability in the target variable that can be explained by the model’s predictors. It can serve as an indicator of the model’s explanatory strength, reflecting how well the observed data points are captured by the regression model.
Unconstrained & Constrained Successful Rate (QED & DRD2). Both unconstrained and constrained success rates calculate the proportion of predicted molecules that satisfy a predefined threshold out of all predicted molecules. The unconstrained variant emphasizes the model’s exploration of a broader chemical space to produce high-quality molecules that meet the specified properties, such as new scaffolds or substituents with improved characteristics. The constrained variant considers molecular similarity and focuses on preserving key molecular features (e.g., scaffolds, pharmacophores). QED and similarity are readily computed by RDKit. However, the commonly used approach for computing the DRD2 score relies on machine learning methods (e.g., support vector machines) for molecular classification [71, 55]. Although research shows that non–deep learning methods generally outperform deep learning approaches on molecular classification tasks [72], this finding may not fully account for the impact of deeper‐level sample distribution factors, such as class imbalance and disparities in feature distributions on dataset size. Moreover, the generalization capability of non–deep learning methods can be limited, which undermines their suitability for DRD2 score computation, since molecules predicted by large language models may be entirely unseen by the DRD2 scoring model. Following [72], we adopt XGBoost [12] as the benchmark representative of machine learning approaches and train GNN-based model GraphMVP [46], on the molecule editing dataset for comparison experimental results and setup details appear in Appendix E. The results indicate that XGBoost easily overfits to the training data, whereas GraphMVP exhibits significantly superior generalization compared to XGBoost, thereby validating the reliability of our proposed scoring approach.
@@ C.1 Graph Tokenizer
Molecule to graph conversion. GNN is widely used in many scenarios, such as traffic [75], social relationships [76], and also molecules [62, 61]. Following the vanilla setting, we utilize RDKit [38] to transform SELFIES into graph structures in our experiments. For tasks involving a single molecule as input, the molecule is converted directly. For tasks requiring multiple molecules as input, only the first molecule in the input sequence is converted into a graph. Our model does not incorporate multi-graph understanding; instead, it processes both the graph and SELFIES representation of the first molecule, while only the SELFIES representations are provided for the remaining molecules. Meanwhile, since MoleculeSTM [45] incorporates additional molecular graph-text contrastive training compared to GraphMVP [46], which leads to improved multimodal model training efficiency, we adopt MoleculeSTM as the graph encoder.
@@
Multiple molecule inputs. In some tasks, (e.g., Reagent Prediction), when we need to copy with multiple molecules as the input, we stack them into batch dimensions and feed them to the graph encoder together, the resulting graph features are correspondingly concatenated together.
@@ C.2 Multimodal Alignment
To balance the molecular graph and text modalities while ensuring training efficiency, we employ a single-layer linear projector in Stage 1. Following [45], we carefully filter PubChem to obtain 310K+ graph-text pairs and convert them into instruction-following data for pretraining. The alignment between the molecular graph and text modalities is enhanced solely by adjusting the parameters of the single-layer linear projector. After that, in the unified instruction tuning stage, we keep the projector active, allowing the projector to adapt to multiple tasks.
C.3 Gradient Adaptive LoRA (GAL)
The scaling factor is calculated as
$\gamma_{\theta}=\frac{\alpha}{r^{p}}+\beta$ (5)
where $\theta=\{\alpha,p,\beta\}$
Inspired by [30, 53, 9, 23], we initialize these parameters as $\alpha_{0}=16,p_{0}=0.5,\beta_{0}=0$. Additionally, we clip the range of these learnable parameters.
$\alpha=\mathrm{clip}(\alpha,\alpha_{0}-\epsilon,\alpha_{0}+\epsilon)\quad p=\mathrm{clip}(p,p_{0}-\delta,p_{0}+\delta)\quad\beta=\mathrm{clip}(\beta,\beta_{0}-\epsilon,\beta_{0}+\epsilon)$ (6)
and we set $\epsilon=0.05$, $\delta=0.01$. If the rank is set to 64, then $\gamma_{\theta}\in[1.863,2.141]$.
C.4 Mutual Representation Similarity
Task scaling setup. We build a sequence of multi-task datasets with detailed composition as follows:
@@ C.4 Mutual Representation Similarity
  • 2 tasks: Reagent Prediction ++ Molecular Captioning.
  • 4 tasks: Reagent Prediction ++ Molecular Captioning ++ Solvent Prediction ++ Catalyst Prediction.
  • 8 tasks: Reagent Prediction ++ Molecular Captioning ++ Solvent Prediction ++ Catalyst Prediction ++ Forward Prediction ++ Retrosynthesis ++ Property Prediction ++ Yield Regression.
@@
where $R^{\prime}\in\mathbb{R}^{B\times L\times d}$, and $m\in\mathbb{R}^{B\times T}$ is the mask indicating the padding tokens. We then flatten the first two dimensions and get $R^{\prime\prime}\in\mathbb{R}^{(B*L)\times d}$ and calculate the similarity with mutual_knn [27].
Let N=BLN=B*L, and we have two models $A$ and $B$ trained on different multi-task datasets, we first find their $k$ nearest neighbors $\text{knn}^{A}$ and $\text{knn}^{B}$.
@@ Appendix D Further details on training
Forward Reaction Prediction 8e-5 15 cosine 10 128 0.0075
@@
Text Guided Molecule Generation
Molecule Editing
IUPAC2SELFIES
@@
Omni-Molecular Tasks (2D & 3D) 15 15
@@
Specialist models are typically fine-tuned on a single task at a time, repeating the process separately for each task, a strategy known as separate tuning. In contrast, generalist models undergo simultaneous fine-tuning across multiple tasks, a process referred to as unified tuning. In this section, we present a detailed training framework for both of them in all experiments.
Separate instruction tuning. We follow the training recipe outlined in [10]. However, we observe significant overfitting when training the model on the molcap task for 20–50 epochs, as suggested in [10]. To address this issue, we manually allocate $10\%$ of the training set for validation and re-evaluated all tasks, we find that the recipes for forward prediction, reagent prediction, retrosynthesis, and Quantum Mechanics Property Prediction from the original paper match our results, however, we identify an updated training strategy tailored to the molcap task. The revised training recipe is summarized in Table 5.
@@
For all experiments, the weight decay is set to 0. The term Stop Epoch in Table 5 shows the epoch that the experiment stops. This is because of the early stop mechanism we used to prevent overfitting.
@@ Appendix E Further details on experimental results
E.1 Baslines
In-Context Learning with DeepSeekV3. To inspect the capability of In-Context Learning with a powerful open-source model like DeepSeekV3, we randomly sample 5 examples from training set and feed the model with instruction, SELFIES representation and the answer. The model is asked to solve the task from test set with these in-context examples. The prompt template is as follows:
Instruction: Given the following examples: {Question: question Answer: answer}*5. Answer the following question, Question: question.
here, question and answer are the specific samples from the dataset, *5 means we provide 5 question-answer pairs.
Molecular LLMs. Mol-Instruction, InstructMol, HIGHT, 3D-MoLM, and PRESTO represent a series of works leveraging large language models (LLMs) to perform molecular tasks. Among them, Mol-Instruction, InstructMol, and HIGHT are specialist models that employ different adapters tailored to specific tasks. 3D-MoLM offers both generalist and specialist variants, and we report the performance of both in Table 3. PRESTO adopts a full fine-tuning strategy using a single model. For datasets and tasks overlapping with ours, we directly report the results from the original paper. For tasks that PRESTO was not designed to handle, we conducted our own re-implementation.
Older Baselines. We include some older baselines such as Llama2 [65](result from [21]), nach0 [49](result from [11]), T5Chem [51](result from [48]), TextChemT5 [14](result from [48]), MolT5 [19](result from [48]), Galatica [64](result from [48]), MolCA [47](result from [48]), GPT-4-0314 [1](result from [10]), BioMedGPT [52](result from [10]).
E.2 3D Adaptive Ability of Omni-Mol
Model #Par Exa BLEU Lev RDK MAC Mor Val
Forward Reaction Prediction Task
Omni-Mol GL 2.2B 0.73 0.980 5.55 0.895 0.947 0.87 1.00
Omni-Mol (3D) GL 2.2B 0.75 0.983 5.56 0.893 0.944 0.88 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
Retrosynthesis Task
Omni-Mol GL 2.2B 0.57 0.960 8.97 0.864 0.909 0.83 1.00
Omni-Mol (3D) GL 2.2B 0.57 0.958 9.24 0.858 0.910 0.82 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
Reagent Prediction Task
Omni-Mol GL 2.2B 0.23 0.726 14.59 0.557 0.627 0.52 1.00
Omni-Mol (3D) GL 2.2B 0.25 0.706 14.35 0.603 0.663 0.56 1.00
Model #Par Exa BLEU Lev RDK MAC Mor Val
Catalyst Prediction
Omni-Mol GL 2.2B 0.73 0.980 5.55 0.895 0.947 0.87 1.00
Omni-Mol (3D) GL 2.2B 0.77 0.836 1.55 0.913 0.898 0.78 1.00
Model Type #Param HOMO LUMO GAP Avg.
Quantum Mechanics Property Prediction Task
Omni-Mol GL 2.2B 0.0038 0.0047 0.0049 0.0044
Omni-Mol (3D) GL 2.2B 0.0032 0.0036 0.0044 0.0037
Model Type #Param B-2 B-4 R-1 R-2 R-L M
Molecular Captioning Task
Omni-Mol GL 2.2B 0.529 0.440 0.604 0.447 0.541 0.571
Omni-Mol (3D) GL 2.2B 0.511 0.421 0.593 0.435 0.531 0.553
Model #Par B-2 B-4 R-1 R-2 R-L M
Description Q&A Task
Omni-Mol GL 2.2B 0.52 0.44 0.53 0.38 0.49 0.58
Omni-Mol (3D) GL 2.2B 0.51 0.44 0.53 0.39 0.49 0.58
Model #Par Weight LogP
More Mol2Num
Omni-Mol GL 2.2B 11.07(100) 0.49(100)
Omni-Mol (3D) GL 2.2B 10.98(100) 0.48(100)
@@ E.1 Baslines
Model #Param Exact BLEU Levenshtein RDK MACCS Morgan Validity
Solvent Prediction
Omni-Mol GL 2.2B 0.52 0.759 2.71 0.671 0.673 0.64 1.00
Omni-Mol (3D) GL 2.2B 0.56 0.77 2.44 0.70 0.70 0.68 1.00
Table 6: Additional results of Omni-Mol with 3D GNN.
In this section, we answer the following question: Is it possible to develop a generalist 3D molecular LLM capable of effectively learning across diverse task domains? 3D molecules represent the physical form encountered in the real world, revealing richer layers of molecular information. Hence, enabling LLMs to achieve strong performance on 3D tasks is critically important. To more effectively encode 3D molecular graph data, we employ Uni-Mol [81] as the graph encoder, integrating it into our graph tokenizer. To obtain the 3D molecular graph information for the 3D Omni-Mol data, we follow the preprocessing pipeline outlined in Appendix B. As the results shown in Table 6, Omni-Mol achieves state-of-the-art performance on all ten 3D tasks, demonstrating its strong adaptability to tasks that more closely reflect real-world scenarios and its potential significance for practical applications. For 3D unified training, we follow the hyperparameters and training strategies of the 2D Omni-Mol unified training. More details are provided in Appendix D.
E.3 Comparison on Mol2Num Tasks with GNN
In this section, we re-implement GraphMVP as our baseline, we re-train GraphMVP on quantum mechanics property prediction, LogP prediction, molecular weight prediction, TPSA prediction tasks.
As shown in Table 8, Omni-Mol performs significantly better than traditional GNN models like GraphMVP. Omni-Mol provides up to $91\%$ improvement on tasks like molecular weight prediction. GraphMVP-G introduces generative 2D self-supervised learning, i.e., it trains the GNN to reconstruct the attributes of nodes and edges that are randomly masked. GraphMVP-C introduces contrastive 2D self-supervised learning, i.e., it trains the GNN to discriminate between constructed positive and negative molecular-graph sample pairs.
E.4 Design Computational Tool for DRD2 Score
   Model    Acc (Training)    Acc (Test)    ROC-AUC (Training)    ROC-AUC (Test)
   Molecule Editing
   XGBoost    0.9995    0.9284    0.9976    0.7933
   GraphMVP    0.8919    0.8748    0.9482    0.9188
Table 7: Comparison between XGBoost and GraphMVP on DRD2 score.
Since existing methods for computing the DRD2 score do not necessarily generalize to data with different distributions [55, 71], in this session we design a new tool for computing the DRD2 score and demonstrate the superiority of the proposed approach through our experiments. For the Molecule Editing task dataset, we extract unique molecules and the DRD2 score labels provided by [71]. To ensure that the DRD2 score computation method generalizes across different distributions, we enforce minimal similarity among molecules in the training, validation, and test sets, and maintain balanced active and inactive samples with the same threshold of activity in [55]. Following [55], we employ the Butina clustering algorithm [8], yielding 50,552, 5,159, and 6,169 samples in the training, validation, and test sets, respectively. As stated in Appendix B, we choose XGBoost and GraphMVP as representative non–deep-learning and deep-learning models, respectively. Adhering to the training setup of [55], the performance is reported in Table 7. These results demonstrate that GraphMVP attains superior generalization on the Molecule Editing data, which underpins the reliability of DRD2 score evaluation for molecule editing tasks, since the generated molecules are unlikely to have been encountered by the evaluation model during training.
Model HOMO LUMO GAP Weight LogP TPSA
Mol2Num
GraphMVP SL 0.0062 0.0082 0.0103 124.08 0.8257 42.20
GraphMVP-G SL 0.3349 0.0076 0.0104 122.88 0.8099 42.33
GraphMVP-C SL 0.0061 0.0076 0.0104 123.57 0.7806 41.79
Omni-Mol GL 0.0038 0.0047 0.0049 11.07 0.49 5.89
Table 8: Comparison between Omni-Mol and GraphMVP